OpenStudy (anonymous):

If f(x)=(x^9−5x√x)/(x) , then f′(x)= ?

OpenStudy (anonymous):

Simplify f(x). Remember, x^2/x can also be written as x^(2-1)

OpenStudy (anonymous):

can u do the first step plz ?

OpenStudy (anonymous):

No sorry you must learn :)

OpenStudy (anonymous):

\[ f(x)=\frac{x^9−5x \sqrt{x}}{x}=\frac{x^9}{x}-\frac{5x\sqrt{x}}{x}\]

OpenStudy (anonymous):

thanks :)